What is 6 divided by 2+3i?
Answer by Alan3354(69443)

What is 6 divided by 2+3i?
----------------
6/(2+3i)
Multiply by the conjugate of the DEN
=6*(2-3i)/(2+3i)*(2-3i)
= (12-18i)/(4+9)
= (12-18i)/13